# W-RAG: Weakly Supervised Dense Retrieval in RAG for Open-domain Question Answering

### v1

@@ Abstract.
In knowledge-intensive tasks such as open-domain question answering (OpenQA), Large Language Models (LLMs) often struggle to generate factual answers relying solely on their internal (parametric) knowledge. To address this limitation, Retrieval-Augmented Generation (RAG) systems enhance LLMs by retrieving relevant information from external sources, thereby positioning the retriever as a pivotal component. Although dense retrieval demonstrates state-of-the-art performance, its training poses challenges due to the scarcity of ground-truth evidence, largely attributed to the high costs of human annotation. In this paper, we propose W-RAG by utilizing the ranking capabilities of LLMs to create weakly labeled data for training dense retrievers. Specifically, we rerank the top-$K$ passages retrieved via BM25 by assessing the probability that LLMs will generate the correct answer based on the question and each passage. The highest-ranking passages are then used as positive training examples for dense retrieval. Our comprehensive experiments across four publicly available OpenQA datasets demonstrate that our approach enhances both retrieval and OpenQA performance compared to baseline models. Source code is published  [^1] https://github.com/jmnian/weak_label_for_rag.
Retrieval Augmented Generation, Open-domain Question Answering, Weak Supervised Learning, Dense Retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Open-domain question answering (OpenQA) dating back to the 1960s (Jr. et al., 1961) provides natural language-like answers to reply users’ questions. OpenQA adopts the “Retriever-Reader” architecture (Zhu et al., 2021), where the retriever retrieves relevant passages for the reader to generate answers. Previous studies (Lewis et al., 2020; Xiong et al., 2020a) adopt a seq2seq model as a reader and train it on the labeled dataset. Recently, large language models (LLMs) like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have demonstrated astonishing performance on various tasks, including OpenQA, attributing to the substantial amount of knowledge stored in their internal parameters. Despite the unprecedented achievements of LLMs, they face constraints such as the inability to consistently integrate up-to-date knowledge, as their parametric knowledge is fixed after being trained on huge datasets. Additionally, they are prone to generating plausible but non-factual responses, known as hallucinations (Welleck et al., 2020).
@@
To overcome the limitations of LLMs’ parametric knowledge, retrieval augmented generation (RAG) (Lewis et al., 2020; Guu et al., 2020) is explored, equipping LLMs with a retriever to gather necessary evidence from external sources. Among the two components of RAG, improving the retriever is more feasible due to the recent trend of black-box APIs (OpenAI, 2023) and the high cost and time requirements of fine-tuning open-source LLMs (Dubey et al., 2024). The retriever, a critical part of RAG, is typically either a traditional unsupervised retriever like BM25 (Robertson and Zaragoza, 2009) or a more advanced neural retriever, such as dense retrieval (Nogueira et al., 2020; Xiong et al., 2020b; Khattab and Zaharia, 2020a; Karpukhin et al., 2020b), which encodes questions and passages into the same embedding space and then measures the question-passage relevance score by vector similarity. A key challenge in training dense retrievers is the scarcity of human-annotated data, as in OpenQA, human-labeled evidence passages are often unavailable. Methods like UPR (Sachan et al., 2022a), which ranks passages based on the likelihood of LLMs generating the question given the passage, and AAR (Yu et al., 2023), which combines the top-$K$ passages ranked by the LLM’s averaged cross-attention scores with ground-truth passages as positive passages, have been employed to train dense retrievers. These approaches can train retrievers to find semantically relevant passages but do not guarantee improved RAG performance in OpenQA. As Cuconasu et al. (2024) demonstrated, retrieving relevant passages that cannot answer the question may negatively impact RAG performance in OpenQA tasks. Neither cross-attention scores nor the likelihood of LLMs generating the question based on the input passage explicitly ensures that the question can be answered by the “positive” passage.
@@
To address the scarcity of training data for dense retrievers in RAG for OpenQA, we propose extracting weak labels from existing OpenQA question-answer pairs by leveraging the ranking capabilities of LLMs (Sun et al., 2023). Specifically, we first use BM25 to retrieve the top-$K$ passages for a question, then pair each passage with the question. We rank the passages based on the likelihood that the LLM would generate the question’s ground-truth answer from each question-passage pair. Only the top-ranked passage is selected as the positive example for the question, and we train the dense retrievers using in-batch negative sampling. Our method evaluates the relevance of the question-passage pair by the likelihood of the answer given a passage and the question, making it particularly suitable for OpenQA, where retrieved passages must be capable of providing the correct answer.
@@
  • We propose W-RAG, a general framework to generate weak labels from question-answer pairs for training dense retrievers in RAG.
  • Comprehensive experiments were conducted, and the results prove that W-RAG improves both retrieval and OpenQA performance over baselines.
  • Open source LLMs were used, and the code was released anonymously to ensure reproducibility.
@@ 2.1. Dense Retrieval
Traditional information retrieval (IR) method are based on exact term matching, like BM25. While it is still widely used due to its efficiency, effectiveness, and robustness, it suffers from the well-known issue of lexical gap (Berger et al., 2000). To address this, leveraging the neural networks, dense retrieval (DR) employs pre-trained language models like BERT (Devlin et al., 2019), to encode questions and passages into embeddings, and measures the question-passage relevance score by the vector similarly in the embedding space. Specifically, DP encodes the whole corpus into embeddings and builds the index, such as Faiss (Douze et al., 2024), on them. When a new question comes in, DR encodes the question into an embedding and performs a nearest neighbor search. DR can be classified into two categories: supervised, like DPR (Karpukhin et al., 2020b), TAS-B (Hofstätter et al., 2021), and ColBERT (Khattab and Zaharia, 2020a; Santhanam et al., 2022); unsupervised, like Contriever(Izacard et al., 2022) and ReContriever(Lei et al., 2023a), based on the training method.
DPR utilizes a dual-tower architecture, with one BERT model dedicated to encode questions and another to encode passages. The similarity between the question and passage embeddings is then calculated, aiming to maximize the log-likelihood of the positive passage. ColBERT uses the same BERT model for both the question and passage encoders, differentiating them by appending a unique special token after the [CLS] token. Unlike DPR, which directly compares question and passage embeddings, ColBERT introduces a late interaction mechanism. It computes the similarity between each question token and all passage tokens, followed by maximum pooling over these similarities. The final similarity score for a question-passage pair is the sum of the pooled scores. TAS-B groups queries based on their embedding similarities and applies a training data sampling technique along with dual-teacher supervision distillation. Contriever trains a bi-encoder model using contrastive learning, generating positive question-passage pairs from an unlabeled corpus. ReContriever follows the same approach as Contriever for generating weak question-passage pairs but adds a self-scoring mechanism during training, where the loss is weighted by these scores.
@@ 2.2. RAG for OpenQA
RAG models have been applied to OpenQA, demonstrating significant performance improvement. Different RAG models have been proposed to solve the critical issues in OpenQA, such as how to retrieve relevant passages (Yu et al., 2023; Shi et al., 2023; Shao et al., 2023; Trivedi et al., 2023), when to call the retriever (Jiang et al., 2023b; Wang et al., 2023; Cheng et al., 2024), and how to decrease the computational complexity (Li et al., 2023; Jiang et al., 2023a; Xu et al., 2024; Kim et al., 2024).
@@
Since the retriever is a critical component in RAG, some studies have tried to improve the quality of retrieved passages, including training a better retriever (Yu et al., 2023; Shi et al., 2023) and prompt engineering (Shao et al., 2023; Trivedi et al., 2023). REPLUG (Shi et al., 2023) trains a dense retriever by minimizing the KL divergence between the retrieval likelihood and the language model likelihood computed as the normalized language model probability of the ground-truth, given the question and passage. AAR (Yu et al., 2023) combines the top-$K$ passages ranked by the LLM’s averaged corss-attention scores with ground-truth passages as positive passages and then follow ANCE (Xiong et al., 2021) to sample negative passages to train the dense retriever. ITER-RETGEN (Shao et al., 2023) leverages the model output from the previous iteration as a specific context to help retrieve more relevant knowledge. IRCoT (Trivedi et al., 2023) also adopt a similar approach to perform retrieval, but applies CoT for generating responses.
@@
In some cases, LLMs can generate factual content well without external knowledge, and retrieving passages will decrease RAG’s performance. Self-RAG (Wang et al., 2023) and FLARE (Jiang et al., 2023b) both fine-tune the LLMs to call the search engine automatically when external knowledge is needed. UAR (Cheng et al., 2024) trains classifiers to identify the need for external knowledge.
@@
Applying all the top-$K$ retrieved passages as context not only increases the computational complexity and inference latency but also brings noise. Numerous methods have been proposed to compress the retrieved passages. Selective-Context (Li et al., 2023) filters non-essential lexical units by the summarization of self-information of each token contained in the unit. LongLLMLingua (Jiang et al., 2023a) contrasts the perplexity score of each token in the passage with the perplexity score of the same token conditioned on the question and adopts this conservative perplexity score to filter out tokens. RECOMP (Xu et al., 2024) leverages the summarization models to summarize the retrieved passages. SuRe (Kim et al., 2024) summarizes retrieved passages conditioned on each answer candidates generated by prompting the LLMs and then selects the top summarizations ranked by LLMs through a combination of pointwise and pairwise scoring method.
@@
Figure 1. W-RAG fits into the general RAG pipeline by training the retriever with LLM generated weak labels. Following the steps from top to bottom, we retrieve candidate passages using BM25, present each passage to the LLM to rerank based on the answer likelihood, then use the reranked top passage to train the retriever to finally enhance LLM’s response quality through the standard RAG pipeline. The entire prompt is shown in Figure 2.
@@ 2.3. LLMs for Ranking
Concerning the generation of weakly labeled ranking data using LLMs as passage rankers, existing work can be categorized into three categories. The first involves a zero-shot listwise ranking strategy, where a prompt containing the question, instructions, and multiple passages are given to the LLM, and the LLM generates a permutation of ranked indices (Baldelli et al., 2024; Ma et al., 2023; Pradeep et al., 2023; Shen et al., 2023). The second category prompts the LLM with passages and directly instructs it to rate their relevance (Zhuang et al., 2024), or perform pairwise comparisons (Qin et al., 2023). The third approach calculates the question generation likelihood when a passage is included in the prompt (Cho et al., 2023; Sachan et al., 2022b; Zhuang et al., 2023). Our method for scoring passages aligns with this third approach; however, as mentioned in Section 1, we rank the passages by the likelihood of the LLM generating the question’s ground-truth answer based on the question-passage pair.
@@ 3. Method
As illustrated in Figure 1, W-RAG introduces a novel approach for training the retrieval component within the RAG pipeline from scratch. Our method leverages a weakly-supervised training paradigm that requires only a set of question-answer pairs and an evidence corpus. The W-RAG process unfolds in three stages: first, it retrieves relevant passages from the evidence corpus; second, it employs a LLM to generate weak labels from them by reranking the retrieved passages by the likelihood of generating the ground-truth answer; finally, the dense retiever is trained using these weak labels. The resulting dense retriever is then capable of retrieving evidence, thereby enhancing the performance of LLMs across a wide array of tasks.
Figure 2. Prompts used for weak label generation and question answering
@@ 3.1. Weak-label Generation
Given an evidence corpus $\mathcal{C}=\{d_{1},\dots,d_{N}\}$, we use a simple retreiver to conduct first-stage retrieval. Given a question $q$, the retriever ranks each passage by relevance and selects the top-$K$ passages $\mathcal{P}=\{s_{1},\dots,s_{K}\}$, where $\mathcal{P}\subset\mathcal{C}$. The corpus may comprise chuncked Wikipedia articles or from online sources such as Reddit and Common Crawl, while the question-answer pairs could be sourced from OpenQA platforms such as Quora and StackOverflow. The simple first-stage retriever could be lexical such as BM25 (Robertson and Zaragoza, 2009), or a dense retriever. The primary objective at this stage is to maximize recall within a manageable set of retrieved passages. Ideally, this step curates evidence passages containing the correct answer to the question
W-RAG is motivated by the hypothesis that an autoregressive LLM is more likely to generate the correct answer if the provided passage contains the necessary information to answer the question. Our approach aims to harness signals from the LLM’s downstream tasks to evaluate the effectiveness of a passage in eliciting the correct answer. It is important to note that traditional relevance measures, such as term overlap or semantic similarity, do no necessarily guarantee that the most relevant passage contains the needed information. Instead, our target is on identifying passages that most effectively prompt the LLM to produce the correct answer and using these passages to train the dense retriever. In this context, we redefine passage relevance as the degree to which a passage can elicit the ground-truth answer, and this criterion serves as the target for training the dense retriever.
@@
As depicted in Figure 1 Step 2, we construct weak label generation prompts with candidate passage $s_{i}$, question $q$, some instructions $I$, and ground-truth answer $a$. The downstream signal is captured through the conditional probability of the ground-truth answer, denoted as $p(a|s_{i},q,I)$:
(1) $p(a|s_{i},q,I)=\prod_{j}\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$
@@
For each token in the input prompt, the LLM assigns logits, which after applying a softmax operation, correspond to the probability of each vocabulary token being the next in the sequence. The probability of the $j$th answer token, denoted as $\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$, is directly extracted from the logits of the preceding token. Since the ground-truth answer typically consists of multiple tokens, the cumulative probability can diminish rapidly. To address this, we simply take the logarithm on both sides of Equation 1, converting each token’s probability into log-likelihood. The overall probability of the answer $a$ is then computed as the average of the sum of the log-likelihoods for each token in the answer:
(2) $\log p(a|s_{i},q,I)=\frac{1}{|a|}\sum_{j}\log\text{LLM}(a_{j}|s_{i},q,I,a_{<j})$
@@
We consider this log-likelihood as the relevance score for the candidate passage $s_{i}$. Each of the $K$ candidate passages is weakly labeled in this manner, resulting in a ranked list of passages, denoted as $\mathcal{S}=\{s_{r_{1}},\dots,s_{r_{K}}\}$. This list is produced by sorting the set of candidate passages $\mathcal{S}$ according to their relevance scores.

### v2

@@ Abstract.
In knowledge-intensive tasks such as open-domain question answering (OpenQA), large language models (LLMs) often struggle to generate factual answers, relying solely on their internal (parametric) knowledge. To address this limitation, Retrieval-Augmented Generation (RAG) systems enhance LLMs by retrieving relevant information from external sources, thereby positioning the retriever as a pivotal component. Although dense retrieval demonstrates state-of-the-art performance, its training poses challenges due to the scarcity of ground-truth evidence, largely attributed to the high costs of human annotation. In this paper, we propose W-RAG, a method that draws weak training signals from the downstream task (such as OpenQA) of an LLM, and fine-tunes the retriever to prioritize passages that most benefit the task. Specifically, we rerank the top-$k$ passages retrieved via BM25 by assessing the probability that the LLM will generate the correct answer for a question given each passage. The highest-ranking passages are then used as positive fine-tuning examples for dense retrieval. We conduct comprehensive experiments across four publicly available OpenQA datasets to demonstrate that our approach enhances both retrieval and OpenQA performance compared to baseline models, achieving results comparable to models fine-tuned with human-labeled data. Source code is published  [^1] https://github.com/jmnian/WRAG.
@@ 1. Introduction
Open-domain question answering (OpenQA) dating back to the 1960s (Jr. et al., 1961) provides natural language-like answers to reply users’ questions. OpenQA adopts the “Retriever-Reader” architecture (Zhu et al., 2021), where the retriever retrieves relevant passages for the reader to generate answers. Previous studies (Lewis et al., 2020; Xiong et al., 2020a) adopt a seq2seq model as a reader and train it on the labeled dataset. Recently, large language models (LLMs) like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have demonstrated astonishing performance on various tasks, attributed to the substantial amount of knowledge stored in their internal parameters. Despite the unprecedented achievements of LLMs, they face constraints such as the inability to consistently integrate up-to-date knowledge, as their parametric knowledge is fixed after being trained on huge datasets. Additionally, they are prone to generating plausible but non-factual responses, known as hallucinations (Welleck et al., 2020).
To overcome the limitations of LLMs’ parametric knowledge, retrieval augmented generation (RAG) (Lewis et al., 2020; Guu et al., 2020) is explored, equipping LLMs with a retriever to gather necessary evidence from external sources. Among the two components of RAG, improving the retriever is more feasible due to the recent trend of black-box APIs (OpenAI, 2023) and the high cost and time requirements of fine-tuning open-source LLMs (Dubey et al., 2024).
@@
The retriever, a critical part of RAG, is typically either a traditional unsupervised retriever like BM25 (Robertson and Zaragoza, 2009) or a more advanced neural retriever, such as dense retrieval (Nogueira et al., 2020; Xiong et al., 2020b; Khattab and Zaharia, 2020a; Karpukhin et al., 2020b), which encodes questions and passages into the same embedding space and then measures the question-passage relevance score by vector similarity. A key challenge in training dense retrievers is the scarcity of human-annotated data, as in OpenQA, human-labeled evidence passages are often unavailable. Methods like UPR (Sachan et al., 2022a), which ranks passages based on the likelihood of LLMs generating the question given the passage, and AAR (Yu et al., 2023), which combines the top-$k$ passages ranked by the LLM’s averaged cross-attention scores with ground-truth passages as positive passages, have been employed to train dense retrievers. These approaches can train retrievers to find semantically relevant passages but do not guarantee improved RAG performance in OpenQA. As Cuconasu et al. (2024) demonstrated, retrieving relevant passages that cannot answer the question may negatively impact RAG performance in OpenQA tasks. Neither cross-attention scores nor the likelihood of LLMs generating the question based on the input passage explicitly ensures that the question can be answered by the “positive” passage.
@@
To address the scarcity of training data for dense retrievers in RAG for OpenQA, and to bridge the gap between semantic relevance and containing the actual answer, we propose extracting weak labels from existing OpenQA question-answer pairs by leveraging the ranking capabilities of LLMs (Sun et al., 2023). Specifically, we first use BM25 to retrieve the top-$k$ passages for a question, then pair each passage with the question. We rank the passages based on the likelihood that the LLM would generate the question’s ground-truth answer from each question-passage pair. Only the top-ranked passage is selected as the positive example for the question, and we train the dense retrievers using in-batch negative sampling. Our method evaluates the relevance of a question-passage pair by the likelihood of generating the correct answer for the question given the passage, making it particularly suitable for OpenQA, where retrieved passages must be capable of providing the correct answer.
@@
  • We propose W-RAG, a general framework that uses answer likelihoods from an LLM given question-answer pairs to generate weak relevance labels for fine-tuning dense retrievers in RAG systems.
  • Comprehensive experiments demonstrate that W-RAG data improves both retrieval and OpenQA performance over baselines and achieves results comparable to models trained on human-labeled data.
  • Open source LLMs were used, and the code was released to ensure reproducibility.
@@ 2.1. Dense Retrieval
Traditional information retrieval (IR) methods are based on exact term matching, like BM25. While it is still widely used due to its efficiency, effectiveness, and robustness, it suffers from the well-known issue of lexical gap (Berger et al., 2000). To address this, leveraging neural networks, dense retrieval (DR) employs pre-trained language models like BERT (Devlin et al., 2019) to encode questions and passages into embeddings, and measures the question-passage relevance score by comparing the vectors similarly in the embedding space. Specifically, DR encodes the whole corpus into embeddings and builds the index, such as Faiss (Douze et al., 2024), on them. When a new question comes in, DR encodes the question into an embedding and performs a nearest neighbor search. DR can be classified into two categories: supervised, like DPR (Karpukhin et al., 2020b), TAS-B (Hofstätter et al., 2021), and ColBERT (Khattab and Zaharia, 2020a; Santhanam et al., 2022); unsupervised, like Contriever(Izacard et al., 2022) and ReContriever(Lei et al., 2023a), based on the training method.
DPR utilizes a dual-tower architecture, with one BERT model dedicated to encoding questions and another to encoding passages, though in practice sometimes a single encoder is shared between passages and questions. The similarity between the question and passage embeddings is calculated, aiming to maximize the log-likelihood of the positive passage. ColBERT uses the same BERT model for both the question and passage encoders, differentiating them by appending a unique special token after the [CLS] token. Unlike DPR, which directly compares question and passage embeddings, ColBERT introduces a late interaction mechanism. It computes the similarity between each question token and all passage tokens, followed by maximum pooling over these similarities. The final similarity score for a question-passage pair is the sum of the pooled scores. TAS-B groups queries based on their embedding similarities and applies a training data sampling technique along with dual-teacher supervision distillation. Contriever trains a bi-encoder model using contrastive learning, generating positive question-passage pairs from an unlabeled corpus. ReContriever follows the same approach as Contriever for generating weak question-passage pairs but adds a self-scoring mechanism during training, where the loss is weighted by these scores.
@@ 2.2. RAG for OpenQA
RAG models have been applied to OpenQA, demonstrating significant performance improvement. Different RAG models have been proposed to solve the critical issues in OpenQA, such as how to retrieve relevant passages (Yu et al., 2023; Shi et al., 2024; Shao et al., 2023; Trivedi et al., 2023), when to call the retriever (Jiang et al., 2023c; Wang et al., 2023; Cheng et al., 2024), and how to decrease the computational complexity (Li et al., 2023; Jiang et al., 2023b; Xu et al., 2024; Kim et al., 2024).
@@
Since the retriever is a critical component in RAG, some studies have tried to improve the quality of retrieved passages, including training a better retriever (Yu et al., 2023; Shi et al., 2024) and prompt engineering (Shao et al., 2023; Trivedi et al., 2023). AAR (Yu et al., 2023) combines the top-$k$ passages ranked by the LLM’s averaged cross-attention scores with ground-truth passages as positive passages and then follows ANCE (Xiong et al., 2021) to sample negative passages to train the dense retriever. ITER-RETGEN (Shao et al., 2023) leverages the model output from the previous iteration as a specific context to help retrieve more relevant knowledge. IRCoT (Trivedi et al., 2023) also adopts a similar approach to perform retrieval, but applies CoT for generating responses.
@@
In some cases, LLMs can generate factual content well without external knowledge, and retrieving passages will decrease RAG’s performance. Self-RAG (Wang et al., 2023) and FLARE (Jiang et al., 2023c) both fine-tune the LLMs to call the search engine automatically when external knowledge is needed. UAR (Cheng et al., 2024) trains classifiers to identify the need for external knowledge.
@@
Applying all the top-$k$ retrieved passages as context not only increases the computational complexity and inference latency but also brings noise. Numerous methods have been proposed to compress the retrieved passages. Selective-Context (Li et al., 2023) filters non-essential lexical units by the summarization of self-information of each token contained in the unit. LongLLMLingua (Jiang et al., 2023b) contrasts the perplexity score of each token in the passage with the perplexity score of the same token conditioned on the question and adopts this conservative perplexity score to filter out tokens. RECOMP (Xu et al., 2024) leverages the summarization models to summarize the retrieved passages. SuRe (Kim et al., 2024) summarizes retrieved passages conditioned on each answer candidate generated by prompting the LLMs and then selects the top summarizations ranked by LLMs through a combination of pointwise and pairwise scoring methods.
@@
Figure 1. W-RAG fits into the general RAG pipeline by training the retriever with LLM-generated weak labels. Following the steps from top to bottom, we retrieve candidate passages using BM25, present each passage to the LLM to rerank based on the answer likelihood, then use the reranked top passage to train the retriever to finally enhance LLM’s response quality through the standard RAG pipeline. Prompt in “Step 2” is shown in Figure 2.
2.3. LLM Signals for Retriever Training
REPLUG (Shi et al., 2024) improves dense retrieval by minimizing the KL divergence between the retrieval likelihood and LM likelihood of generating the second half of a text given the first half and a retrieved passage. The higher the likelihood that the LM will generate the second half correctly, the more relevant the retrieved passage is considered. Since only two separate corpora of text are needed to mine training signals, the method is highly scalable. CoRAG (Wang et al., 2025) while focusing on multi-hop retrieval, selects training samples based on retrieval chains that maximize answer likelihood from the LLM. (Li et al., 2024) propose a teacher-student distillation method where they prompt the LLM to directly output ranked lists to fine-tune a reranker via list-wise ranking loss, while simultaneously aligning the retriever’s output distribution with the reranker’s. Similarly, ARL2 (Zhang et al., 2024) scores relevance by prompting LLMs to annotate passages as either supporting or not supporting the answer for a given question, then trains the retriever to align with these annotations.
@@ 2.4. LLMs for Ranking
Concerning the generation of weakly labeled ranking data using LLMs as passage rankers, existing work can be broadly grouped into three categories. The first category involves a zero-shot listwise ranking strategy, where a prompt containing the question, instructions, and multiple passages is given to the LLM, and the LLM generates a permutation of ranked indices (Baldelli et al., 2024; Ma et al., 2023; Pradeep et al., 2023; Shen et al., 2023). The second category focuses on pointwise and pairwise relevance assessment. (Zhuang et al., 2024) prompts the LLM with a passage and directly instructs it to rate its relevance, making it a point-wise relevance assessment method. (Qin et al., 2023) performs pair-wise ranking where two passages are given in the prompt and the LLM is asked to point out which is more relevant to the question. The third approach calculates the question generation likelihood when a passage is included in the prompt (Cho et al., 2023; Sachan et al., 2022b; Zhuang et al., 2023). Our method for scoring passages aligns with this third approach. However, as mentioned in the introduction, our method ranks passages based on the likelihood that the LLM generates the ground-truth answer when conditioned on the question-passage pair. The ranked lists are then used as weakly labeled training data to fine-tune the retriever, with the goal of teaching it to retrieve passages that are most likely to condition the LLM to generate a correct answer.
Prompt for Answer Likelihood Extraction Passage: {passage} Question: {question} Please answer the question using the facts of the passage. Keep your answer grounded to the facts of the passage. Keep your answer within one short sentence. Answer: {answer}
Figure 2. Bolded portion is evaluated for answer likelihood, which serves as a weak relevance score between “passage” and “question”.
@@ 3. Method
As illustrated in Figure 1, W-RAG introduces a novel approach for fine-tuning the retrieval component within the RAG pipeline from a pre-trained dense retriever. Our method employs a weakly supervised training paradigm that requires only a set of question-answer pairs and a passage corpus. The W-RAG process unfolds in three stages: First, a first-stage retriever, such as BM25, retrieves candidate passages from the corpus. Second, we use an LLM to estimate the relevance of each passage by computing the likelihood of generating the correct answer when conditioned on the given question and passage, producing a weakly labeled ranking list. Finally, the dense retriever is fine-tuned using these weak labels. The resulting dense retriever is optimized for ranking passages with higher answer likelihood towards top positions, thereby enhancing the performance of LLMs across a wide array of tasks.
@@ 3.1. Weak-label Generation
Given a passage corpus $\mathcal{C}=\{s_{1},\dots,s_{N}\}$, we use a simple retriever to conduct first-stage retrieval. Given a question $q$, the retriever ranks each passage by relevance and selects the top-$k$ passages $\mathcal{P}=\{s^{\prime}_{1},\dots,s^{\prime}_{K}\}$, where $\mathcal{P}\subset\mathcal{C}$. The passage corpus may comprise chunked Wikipedia articles or those from online sources, such as Reddit and Common Crawl, while the question-answer pairs could be sourced from QA datasets and question-answering platforms, such as Quora and Stack Overflow. The simple first-stage retriever could be lexical, such as BM25 (Robertson and Zaragoza, 2009), or a dense retriever. The primary objective at this stage is to maximize recall within a manageable set of retrieved passages. Ideally, this step fetches evidence passages containing the correct answer to the question.
W-RAG is based on the hypothesis that an autoregressive LLM is more likely to generate the correct answer when the input passage contains the necessary information. We aim to leverage this signal by fine-tuning the retriever to rank passages according to their likelihood of eliciting the correct answer. Unlike traditional relevance measures, such as term overlap or semantic similarity, which do not emphasize that a passage enables answer generation. Our approach redefines relevance as the extent to which a passage conditions the LLM to produce the ground-truth answer. These labels serve as a weak supervision signal for fine-tuning the retriever.
@@
As depicted in Figure 1 Step 2, we construct weak label generation prompts with candidate passage $s^{\prime}_{i}$, question $q$, some instructions $I$, and ground-truth answer $a$. The downstream signal is captured through the conditional probability of the ground-truth answer, denoted as $p(a|s^{\prime}_{i},q,I)$:
(1) $p(a|s^{\prime}_{i},q,I)=\prod_{j}\text{LLM}(a_{j}|s^{\prime}_{i},q,I,a_{<j})$
@@
For each token in the input prompt, the LLM assigns logits, which, after applying a softmax operation, correspond to the probability of each vocabulary token being the next in the sequence. The probability of the $j$th answer token, denoted as $\text{LLM}(a_{j}|s^{\prime}_{i},q,I,a_{<j})$, is directly extracted from the logits of the preceding token. Since the ground-truth answer typically consists of multiple tokens, the cumulative probability can diminish rapidly. To address this, we take the logarithm on both sides of Equation 1, converting each token’s probability into log-likelihood. The overall probability of the answer $a$ is then computed as the average of the sum of the log-likelihoods for each token in the answer:
(2) $\log p(a|s^{\prime}_{i},q,I)=\frac{1}{|a|}\sum_{j}\log\text{LLM}(a_{j}|s^{\prime}_{i},q,I,a_{<j})$
@@
We consider this log-likelihood as the relevance score for the candidate passage $s^{\prime}_{i}$. Each of the $k$ candidate passages is weakly labeled in this manner, resulting in a ranked list of passages, denoted as $\mathcal{S}=\{s^{\prime}_{r_{1}},\dots,s^{\prime}_{r_{K}}\}$. This list is produced by sorting the set of candidate passages $\mathcal{S}$ according to their relevance scores.